A right triangle has a leg measuring 52 m and a hypotenuse measuring 63 m. Find the length of the other leg.
Answer by ewatrrr(24785)


Hi,
Let x represent the length of the other leg
Applying the Pythagorean Theorem
52^2 + x^2 = 63^2
x^2 = 63^2 - 52^2
x^2 = 1265m^2
x = 35.567m